Problem 2. Determine the heat of reaction at 25°C for the following elementary chemical reactions and state whether each onePlease help with each reaction!!
Answer #1

02 + N - NO + N Heat of reaction -AH = { AHA (products) - & AM Creactant) 467 X 4+18 (No) + 2H1 CM) – 1H (02) - 444 () + 40N o ( 249 kJ/mol -ve, 06 (-230 kJ OH H01.241 KT.). - AH Со t pH -- Co, f н AhrenAH1(04) + Ath() – 4:14.660) -4210H) = -393 +34. coton - Co to. AMin = 15$ (Co) + 4H.CO) - .(CO) -47.603) -393 - 6419 +249 -(-110) - 0 =-34 kJ. ( 4HCO, Exothermic) 0 (244the to ki to (oks) To (249 kJ ) M(-218 km)